OpenStudy (squirrels):

Can someone explain how to do this problem? Please c:

OpenStudy (squirrels):

How do the graphs of y= (1/x) and y= (5/x+6) compare?

OpenStudy (anonymous):

do u need the whole slope thing??

OpenStudy (squirrels):

I dont know, that was the whole question :/

OpenStudy (anonymous):

Yah idk, im bad with math :P

OpenStudy (anonymous):

1/x is red, 5/x + 6, assuming that's what you meant is green

OpenStudy (nincompoop):

first, compare the graph between 1/x and 5/x then 5/x and 5/x+6 finally, 1/x and 5/x+6 there are shifts and dilation or shrinkage

OpenStudy (squirrels):

and how do you graph them? sorry :p

OpenStudy (nincompoop):

you do create a table x | y = 1/x -3 -1/3 -2 -1/2 -1 0 .... so on and so forth then you plot the points

OpenStudy (squirrels):

i kinda get what your saying, so the line never stops going then?

OpenStudy (nincompoop):

whatever will satisfy a good graph

OpenStudy (nincompoop):

or you can just use wolfram alpha or google graph you can see the trends without plotting points

OpenStudy (squirrels):

so how do they compare then?

OpenStudy (nincompoop):

are you trolling or what? you look at the graph and observe the changes

OpenStudy (squirrels):

lol no im not trolling. Im just that stupid. thanks for pointing that out.

OpenStudy (anonymous):

use wolframalpha.com site and type your equations to see the graphs. For example, y = 1/x , y = 5/(6+x)

OpenStudy (squirrels):

Thank you c:

OpenStudy (anonymous):

I did send you them..

OpenStudy (nincompoop):

he's trolling, he's doing calc 3 and differential equations already

OpenStudy (squirrels):

i swear im not trolling nin. and im doing algebra 2 -.- and oh i didnt see that thanks dc!